OpenStudy (nsh4267):

Write the expression in the form a+bi. 3/2+4i

OpenStudy (anonymous):

I assume that's \[ \frac{3}{2+4i} \] right? Gotta use parentheses.

OpenStudy (nsh4267):

yes sorry.

OpenStudy (anonymous):

multiply numerator and denumerator by 2-4i

OpenStudy (anonymous):

"complex conjugate"

OpenStudy (anonymous):

this is the way to convert it to the requested form .. what will you get if you do so ?

OpenStudy (nsh4267):

(3/10)-(3/5)i

OpenStudy (anonymous):

You want to multiply the equation by 1 by using the conjugate which is \[2-4i\] so multiply by \[\frac{2-4i}{2-4i}\] what this does is gets rid of the i ont he bottom and gets you c-di^2 which will end up being a number on the bottom since \[i^2=-1\] for example if you multiply what your denominator should loo like now you get \[\frac{3(2-4i)}{(2+4i)(2-4i)}=\frac{3(2-4i)}{4-8i+8i-16i^2}=\frac{3(2-4i)}{4-16(-1)}=\frac{3(2-4i)}{4+16}\]

OpenStudy (anonymous):

correct @nsh4267

OpenStudy (anonymous):

(3/10)-(3/5)i -> its good :)